Determine the circumference of the second Bohr orbit of the Hydrogen atom. Use this to
determine the wavelength of the electron in this orbit; the electron's wave must consist of
an integral number of wavelengths about its orbit's circumference (Why?).
orbit circumference = n wavelength

Finally, determine the velocity v of the electron in this orbit using de Broglie's prescription
for the wavelength of matter waves.
wavelength = h / mv

What percentage of the speed of light is this velocity?
Answer #1

Radius of second orbit of hydrogen atom = 0.529*n2 / Z A°

= 0.529×22 = 2.116×10-10 m

The circumference = 2 πR = 2 × 3.14 × 2.116×10-10 m = 13.295×10-10 m =1.33×10-9 m (Answer)

Circumference = n × \lambda

\lambda = circumference /2 = 6.6476×10-10 m = 6.65×10-10 m. (Answer)

\lambda = h/mv

Velocity of electron = h/m\lambda = 6.626×10-34/(9.11×10-31 × 6.65×10-10 ) = 1.094×106 m/s = 1.09×106 m/s. (Answer)

% of velocity as speed of light =( 1.09×106/3×108)×100 %

= 0.3647%

= 0.36 %. (Answer)